OpenStudy (anonymous):

prove the identity cos(3x)=cos^3(x)-3sin^2(x)cos(x), step by step

OpenStudy (raden):

use the formula : cos(A+B) = cosAcosB - sinAsinB so, cos(3x) = cos(x+2x) = ....

OpenStudy (anonymous):

cos(x)*cos(2x)-sin(x)*sin(2x)

OpenStudy (raden):

yes, so far is good. now, use the identity that cos2x = 2cos^2 (x) - 1 and sin(2x) = 2sin(x)cos(x) apply of them to did u get above

OpenStudy (anonymous):

cos(x)*2cos^2(x)-1-sin(x)*2sin(x)cos(x)

OpenStudy (raden):

then simplify it ... cos(x)cos^2 (x) = .... sin(x)*2sin(x)cos(x) = ...

OpenStudy (anonymous):

cos^3(x)-1-3sin^2(x)cos(x)

OpenStudy (anonymous):

what about the 1?

OpenStudy (anonymous):

If you know Euler's formula:\[e^{i\theta}=\cos(\theta)+i\sin(\theta)\]there's another way to do this.

OpenStudy (anonymous):

i need to make a proof, and dont know that formula.

OpenStudy (anonymous):

You don't get the 1. You forgot to multiply the one with cosx

OpenStudy (anonymous):

\[\cos(x)[\cos^{2}(x)-1]-\sin(x)[2\sin(x)\cos(x)]\]

OpenStudy (anonymous):

\[2cos^2(x)\] Made a mistake in the brackets of the first term.

OpenStudy (anonymous):

im not sure what im missing, but what cancels the 1?

OpenStudy (anonymous):

@comput313 I just told you. Why don't you understand what I wrote.

OpenStudy (anonymous):

expand this part first. Tell me what you get. \[cosx(2cos^{2}x-1)\]

OpenStudy (anonymous):

After you expanded that. Expand this: \[sinx(2sinxcosx)=?\]

OpenStudy (anonymous):

cal me an idiot, but i still dont see what removes the 1. cos(x)*cos^2(x)-1=cos^3(x)-1

OpenStudy (anonymous):

You need to learn how to expand properly.

OpenStudy (anonymous):

expand this. x(x+1)=

OpenStudy (anonymous):

what you did above was this. x(x+1)=x^2+1. That's not how you expand.

OpenStudy (anonymous):

This is how you expand. x(x+1)=x^2+x

OpenStudy (anonymous):

i see now, cos(x)*cos^2(x)-1= 2cos^3(x)- 1cos(x) =cos^3(x)

OpenStudy (anonymous):

Now. We're not finished yet.

OpenStudy (anonymous):

factorise the cosx out.

OpenStudy (anonymous):

And that's incorrect.

OpenStudy (anonymous):

2cos^3(x)- 1cos(x) does not equal cos^3(x)

OpenStudy (anonymous):

@comput313

OpenStudy (anonymous):

this is why i have a bad grade in math. please continue explaining.

OpenStudy (anonymous):

Factorise the cosx out. \[2\cos^{3}x-cosx-2\sin^{2}xcosx=?\]

OpenStudy (anonymous):

\[=cosx[2\cos^{2}x-1-2\sin^{2}x]\] You get that right?

OpenStudy (anonymous):

Now You must know this trig equation. \[\sin^{2}x+\cos^{2}x=1\]

OpenStudy (anonymous):

no, let me work through it again

OpenStudy (anonymous):

yes i know that

OpenStudy (anonymous):

Fine you use the double angle results. \[\cos3x=\cos(x+2x)\] \[\cos(x+2x)=\cos(x)\cos2x-\sin(x)\sin2x\]

OpenStudy (anonymous):

no, i got that part

OpenStudy (anonymous):

i factored wrong i got it now

OpenStudy (anonymous):

\[\cos(x)\cos2x-\sin(x)\sin2x=\cos(x)(\cos^{2}x-sin^{2}x)-\sin(x)(2\sin(x)\cos(x))\]

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

We go back to using that trig equation. \[\sin^{2}x+\cos^{2}x=1\]

OpenStudy (anonymous):

So everytime we see 1, we substitute \[\sin^{2}x+\cos^{2}x\]

OpenStudy (anonymous):

ah, i see now

OpenStudy (anonymous):

\[\cos(x)(2\cos^{2}x-1-2\sin^{2}x)\] \[=\cos(x)(2\cos^{2}x-(\sin^{2}x+\cos^{2}x)-2\sin^{2}x)\]

OpenStudy (anonymous):

Now collect like terms and expand again. Once you do that, it will equal the RHS(Right Hand Side) and thus you proved it.

OpenStudy (raden):

well, some formulas i gave above doesnt work... try use the identity cos2x = cos^2 x - sin^2 x, like aztech said let's try again... cos(x+2x) = cosxcos2x - sinxsin2x cos3x = cosx(cos^2 x -sin^2 x) - sinx(2sinxcosx) cos3x = cos^3 x - sin^2 x cosx - 2sin^2 x cosx cos3x = cos^3 x - 3sin^2 x cosx proof

OpenStudy (anonymous):

remember \[-(x+y)=-x-y\] \[-(x+y)\neq -x+y\] I see many students try and do that.

OpenStudy (anonymous):

ok, thank you both

OpenStudy (raden):

yw